OpenStudy (anonymous):

Solve for m: 235 - 8m = 75

OpenStudy (anonymous):

Start with: 235-8m=75 -8m= -160<--------Subtract 235 from both sides m= 20<--------Divide by -8 on both sides

OpenStudy (anonymous):

thanks :)